OpenStudy (anonymous):

System of Equations? 4x - y = 8 -8x + 2y = -5

OpenStudy (anonymous):

Please don't just give me the answer, please just walk me through the steps, I'm pretty sure I'm just messing up a little.

ganeshie8 (ganeshie8):

multiply first equation by -2, what do you get ?

OpenStudy (anonymous):

Sorry, had to go to the bathroom, didn't think I would get help so fast. Would it be x-2y=8?

OpenStudy (anonymous):

Wait nvm u said -2, let me redo it haha

OpenStudy (anonymous):

@ganeshie8 So it would be either x-2y=-8 or x-y=-4. I didin't know if I was suppose to leave it as 2y or what. Or I could be totally wrong and I wasn't even suppose to multiply y by -2 to begin with.

ganeshie8 (ganeshie8):

4x - y = 8 multiply both sides by -2, you would get : -2(4x - y) = -2(8) -8x + 2y = -16

ganeshie8 (ganeshie8):

however second equation says that has to equal -5 : -8x + 2y = -5

ganeshie8 (ganeshie8):

since -5 is not same as -16, the system has no solution.

ganeshie8 (ganeshie8):

graphically, the system represents a pair of parallel lines that never meet

OpenStudy (anonymous):

Okay just to make sure, I was suppose to leave it at -8x+2y=-16. I wasn't suppose to try and solve it haha, don't know why I didnt just do as you said. Do you know a website that I can do a few system of equations with the different methods to practice?